OpenStudy (anonymous):

How would you format sin^2x+cos^2x=1 to show for sinx?

OpenStudy (anonymous):

I don't know how to format \[\sin^2x+\cos^2x=1\] to sin, I think it should be \[sinx=\sqrt{-\cos^2x+1}\]

OpenStudy (anonymous):

But you can't take the square root of a negative number, which is what has me confused.

OpenStudy (jdoe0001):

\(\bf sin^2(\theta)+cos^2(\theta)=1\implies sin^2(\theta)=1-cos^2(\theta) \\ \quad \\ sin(\theta)=\pm\sqrt{1-cos^2(\theta)}\)

OpenStudy (anonymous):

Any thoughts guys?

OpenStudy (anonymous):

So I was partially right?

OpenStudy (jdoe0001):

well.. you're correct.... I simply put the negative cosine at the end, is all

OpenStudy (anonymous):

Okay thanks jdoe!

OpenStudy (jdoe0001):

yw